# The genealogy of nearly critical branching processes in varying environment

### v3

@@ 3.1.2 Truncation in the Gromov-weak topology
Applying the method of moments requires to carry out a truncation of the reproduction laws for all moments to be finite. In order to deduce from the convergence of the truncated BPVE that of the original BPVE, we need to show that the two trees remain close with large probability for an adequate choice of distance between trees. It turns out that one can define a distance on the space of metric measure spaces that induces the Gromov-weak topology [19, 17, 22]. There are several formulations for such a distance, we work with the Gromov–Prohorov distance of [19, 9]. It is defined as
@@
where the infimum is taken over all complete separable metric spaces $(Z,d_{Z})$ and isometric embeddings $\phi\colon X\hookrightarrow Z$ and $\phi^{\prime}\colon X^{\prime}\hookrightarrow Z$, $\nu\circ\phi^{-1}$ is the pushforward measure of $\nu$ through the map $\phi$, and $d_{\mathrm{P}}$ is the Prohorov distance between finite measures on $(Z,d_{Z})$. The following simple lemma will be the key to control the distance between the truncated and the original tree.
@@
Consider a metric measure space $[X,d,\nu]$. Let $X^{\prime}$ be a closed subset of $X$, and let $d^{\prime}$ and $\nu^{\prime}$ be the restrictions of $d$ and $\nu$ to $X^{\prime}$. Then
@@
    we have $[X^{\prime}_{n},d^{\prime}_{n},\nu^{\prime}_{n}]\to[X,d,\nu]$ in distribution for the Gromov-weak topology.
Then $([X_{n},d_{n},\nu_{n}])_{n}$ also converges to $[X,d,\nu]$ in distribution for the Gromov-weak topology.
@@ 3.2 The Gromov–Hausdorff–Prohorov topology
We now turn to the definition of two very related topologies, the Gromov–Hausdorff–Prohorov and the Gromov–Hausdorff-weak topologies. Following [3], we say that a sequence of metric measure spaces $\big{(}[X_{n},d_{n},\nu_{n}]\big{)}_{n\geq 1}$ converges in the Gromov–Hausdorff-weak sense to $[X,d,\nu]$ if there exists a complete separable metric space $(Z,d_{Z})$ and isometries $\phi_{n}\colon\operatorname{supp}\nu_{n}\hookrightarrow Z$ and $\phi\colon\operatorname{supp}\nu\hookrightarrow Z$ such that
@@
Clearly, a sequence of metric measure spaces converges Gromov–Hausdorff-weakly if and only if the sequence of supports of the measures converges in the Gromov–Hausdorff–Prohorov sense. The two notions of convergence coincide when the spaces have full support, that is, when $\operatorname{supp}\nu=X$.
The Gromov–Hausdorff–Prohorov topology is fairly standard and has been used extensively for deriving scaling limits of trees and other metric spaces, which makes it a natural choice for formulating our main result. However, when trying to reinforce a Gromov-weak convergence to a GHP one, one runs into the problem that the two topologies are not formally defined on the same state space. The GHP topology is defined on the space of strong equivalence classes of metric measure spaces, where two such spaces are said to be strongly equivalent if there exists a bijective isometry between them that preserves the measures, see Definition 2.4 in [1]. On the other hand, the Gromov-weak topology is defined on weak equivalence classes of metric measure spaces, where two such spaces are weakly equivalent if their supports are strongly equivalent in the above sense, see Definition 2.1 in [19]. The Gromov–Hausdorff-weak topology is an adaptation of the GHP topology to weak equivalence classes; it is less standard, but more convenient to use when working with moments and the Gromov-weak convergence. In order to obtain a GHP convergence, we will first prove the Gromov-weak convergence, then reinforce it to a Gromov–Hausdorff-weak one, and finally use the fact that the metric measure spaces considered here have full support to deduce the GHP convergence.
@@ 3.2.1 Ultrametric spaces and reduced processes
In this section we construct the reduced process associated to a general ultrametric space and prove that the function mapping an ultrametric space to its reduced process is continuous in the GHP topology.
@@
Suppose that $\big{(}[U_{n},d_{n},\nu_{n}],\,n\geq 1\big{)}$ is a sequence of compact ultrametric measure spaces converging in the Gromov–Hausdorff–Prohorov topology to $[U,d,\nu]$. Then
@@
This shows that there at least $L_{t}(U)$ open balls of $(U_{n},d_{n})$ of radius $t-2\varepsilon$, that is,
@@
If the process $(L_{t}(U),t>0)$ only makes jump of size $1$, the Skorohod convergence follows for instance from [28, Chapter VI, Theorem 2.15]. ∎
@@ 3.2.2 From Gromov-weak to GHP convergence
Let $\big{(}[U_{n},d_{n},\nu_{n}],\,n\geq 1\big{)}$ be a sequence of compact ultrametric measure spaces such that $\operatorname{supp}\nu_{n}=U_{n}$. Suppose that it converges to a limit $[U,d,\nu]$ in distribution in the Gromov-weak topology. If, for all $t>0$, the collection of random variables
@@
is tight, then the sequence also converges in distribution in the Gromov–Hausdorff–Prohorov topology.
@@
In this result, $([U_{n},d_{n},\nu_{n}])_{n}$ should be interpreted as a random sequence of strong isometry classes such that the corresponding sequence of weak isometry classes converges in distribution in the Gromov-weak topology to $[U,d,\nu]$. The limit of $([U_{n},d_{n},\nu_{n}])_{n}$ in the GHP topology is the unique strong isometry class with full support whose weak isometry class is $[U,d,\nu]$.
@@
in the GHP topology. Since $[U_{n},d_{n},\nu_{n}]$ has full support, $\iota([U_{n},d_{n},\nu_{n}])=[U_{n},d_{n},\nu_{n}]$ and the result is proved. ∎
@@ 4 The environmental coalescent point process
Fix some height $t>0$ and consider a Poisson point process $P$ on $[0,\infty)\times(0,t]$ with intensity measure $ds\otimes\nu(dx)$, for some measure $\nu$ on $(0,t]$ such that
@@
see Figure 1 for an illustration. Consider a positive real $m$ and an independent exponentially distributed random variable $Z$ with mean $m$. The CPP at height $t$ associated to $\nu$ and $m$ is the random metric measure space $[(0,Z),d_{\nu},\operatorname{Leb}]$. Its distribution will be denoted by $\operatorname{CPP}_{t}(\nu,m)$. Note that the Brownian CPP of Section 2.3 has distribution $\operatorname{CPP}_{1}(\tfrac{dx}{x^{2}},1)$.
@@
In the definition of a metric measure space of Section 3.1.1, we require the metric space to be complete and the measure to be defined on the corresponding Borel $\sigma$-field. Almost surely $[(0,Z),d_{\nu}]$ is not complete, and we always work with its completion without further mention. The completion can be obtained explicitly by adding a countable number of points as described in [42]. Moreover, the Lebesgue measure $\operatorname{Leb}$ is formally defined on the usual $\sigma$-field on the interval $(0,Z)$, which might be different from the Borel $\sigma$-field of $[(0,Z),d_{\nu}]$. The two $\sigma$-fields actually coincide, as can be seen directly by noting that the balls of $d_{\nu}$ are intervals of $(0,Z)$, see [14, Lemma 1.4] for a formal proof.
For a nearly critical branching process in varying environment, the rescaled genealogy at time $tN$ will have $\operatorname{CPP}_{t}(\nu_{e},\bar{\rho}_{t})$ as its limiting distribution, with $\bar{\rho}_{t}$ as in (16) and
@@
Let $[(0,Z_{B}),d_{B},\operatorname{Leb}]$ be the Brownian CPP with distribution $\operatorname{CPP}_{1}\left(\tfrac{dx}{x^{2}},1\right)$. Recall the expression of $F$,
@@
and consider $F^{-1}$ its right-continuous inverse. Then $[(0,Z_{B}),F^{-1}(d_{B}),\bar{\rho}_{t}\operatorname{Leb}]$ is distributed as $\operatorname{CPP}_{t}(\nu_{e},\bar{\rho}_{t})$, with $\nu_{e}$ defined in (30).
@@
Thus, $[(0,Z_{B}),F^{-1}(d_{B}),\bar{\rho}_{t}\mbox{Leb}]$ is the CPP at height $t$ associated to the measure $\bar{\rho}_{t}\nu_{e}$ and to the random population size $Z_{B}$. By a simple scaling argument, it is not hard to see that the distribution of $[(0,Z_{B}),F^{-1}(d_{B}),\bar{\rho}_{t}\operatorname{Leb}]$ is the same as $[(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]$, the CPP constructed from $\nu_{e}$ and $Z_{e}$. (Note that the total mass needs to be scaled so that the isometry between $[(0,Z_{B}),F^{-1}(d_{B})]$ and $[(0,Z_{e}),d_{\nu_{e}}]$ is measure-preserving.) ∎
@@ 4.1 Moments of the CPP
In order to prove convergence in the Gromov-weak topology we have to know the moments of the environmental CPP. This is provided by the following result.
@@
Let $[(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]$ be the environmental CPP at time $t$. Then for any bounded function $\phi$ with associated polynomial $\Phi$ we have
@@
where $\sigma$ is an independent and uniform permutation of $[k]$ and $\left((H_{i,j})_{i,j}\right)$ has distribution (18).
@@
$\forall s\leq t,\quad\mathbb{P}(H\leq s)=\frac{1/\bar{\rho}_{t}}{\nu_{e}((s,t])+1/\bar{\rho}_{t}}$
holds. This follows from the simple computation
$(\bar{\rho}_{t}\nu_{e}((s,t])+1)^{-1}=\frac{1}{2\bar{\rho}_{t}}\int_{[t-s,t]}e^{X_{t}-X_{u}}\sigma^{2}(du).$
@@ 4.2 Sampling from the CPP
This problem would be much easier if the points were sampled according to an independent Poisson point process with intensity $\theta/\bar{\rho}_{t}dt$. The atoms of such a point process split $(0,Z_{e})$ into a geometric number of subintervals with success parameter $1/(1+\theta)$, and the lengths of these intervals are independent exponential random variables with mean $\bar{\rho}_{t}/(1+\theta)$. Thus, the distances between any two consecutive atoms are independent, and distributed as the maximum of $P$ over an interval with exponential length of mean $\bar{\rho}_{t}/(1+\theta)$. If $H^{\theta}$ denotes a random variable with the same distribution as this maximum, a direct computation shows that
@@
For any $k\geq 1$, let $(U_{1},\dots,U_{k})$ be uniformly distributed on $(0,Z_{e})$, then
[…]
@@
[…]
@@
[…]
@@
Let $[(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]$ have distribution $\operatorname{CPP}_{t}(\nu_{e},\bar{\rho}_{t})$, let $(V_{1},\dots,V_{k})$ be i.i.d. uniform random variables on $(0,1)$ and define $U_{i}=V_{i}Z_{e}$, $i\leq k$. Since $Z_{e}$, $d_{\nu_{e}}$, and $(V_{i})$ are independent, using Lemma 4.5 we can write
[…]
@@
According to the previous discussion, the random variables $(H^{\theta}_{1},\dots,H^{\theta}_{k-1})$ are i.i.d. and distributed as the maximum of the point process $P$ over an interval with exponential length and mean $\bar{\rho}_{t}/(1+\theta)$. The distribution of this maximum is given by (32). By construction of the CPP distance,

### v4

@@ 3.1.2 Truncation in the Gromov-weak topology
Applying the method of moments requires us to carry out a truncation of the reproduction laws for all moments to be finite. In order to deduce the convergence of the original BPVE from that of its truncation, we need to show that the two trees remain close with high probability for an adequate choice of distance between trees. It turns out that one can define a distance on the space of metric measure spaces that induces the Gromov-weak topology [19, 17, 22]. There are several formulations for such a distance, we work with the Gromov–Prohorov distance of [19, 9]. It is defined as
@@
where the infimum is taken over all complete separable metric spaces $(Z,d_{Z})$ and isometric embeddings $\phi\colon X\hookrightarrow Z$ and $\phi^{\prime}\colon X^{\prime}\hookrightarrow Z$, and where $\nu\circ\phi^{-1}$ is the pushforward measure of $\nu$ by the map $\phi$, and $d_{\mathrm{P}}$ is the Prohorov distance between finite measures on $(Z,d_{Z})$. The following simple lemma will be the key to control the distance between the truncated and the original tree.
@@
Consider a metric measure space $X,d,\nu]$. Let $X^{\prime}$ be a closed subset of $X$, and let $d^{\prime}$ and $\nu^{\prime}$ be the restrictions of $d$ and $\nu$ to $X^{\prime}$. Then
@@
    we have $X^{\prime}_{n},d^{\prime}_{n},\nu^{\prime}_{n}]\to[X,d,\nu]$ in distribution for the Gromov-weak topology.
Then $([X_{n},d_{n},\nu_{n}])_{n}$ also converges to $X,d,\nu]$ in distribution for the Gromov-weak topology.
@@ 3.2 The Gromov–Hausdorff–Prohorov topology
We now turn to the definition of two very related topologies, the Gromov–Hausdorff–Prohorov and the Gromov–Hausdorff-weak topologies. Following [3], we say that a sequence of metric measure spaces $\big([X_{n},d_{n},\nu_{n}]\big)_{n\geq 1}$ converges in the Gromov–Hausdorff-weak sense to $X,d,\nu]$ if there exists a complete separable metric space $(Z,d_{Z})$ and isometries $\phi_{n}\colon\operatorname{supp}\nu_{n}\hookrightarrow Z$ and $\phi\colon\operatorname{supp}\nu\hookrightarrow Z$ such that
@@
Clearly, a sequence of metric measure spaces converges Gromov–Hausdorff-weakly if and only if the sequence of the supports of the measures converges in the Gromov–Hausdorff–Prohorov sense. The two notions of convergence coincide when the spaces have full support, that is, when $\operatorname{supp}\nu=X$.
The Gromov–Hausdorff–Prohorov topology is fairly standard and has been used extensively for deriving scaling limits of trees and other metric spaces, which makes it a natural choice for formulating our main result. However, when trying to reinforce a Gromov-weak convergence to a GHP one, one runs into the problem that the two topologies are not formally defined on the same state space. The GHP topology is defined on the space of strong equivalence classes of metric measure spaces: two spaces are said to be strongly equivalent if there exists a bijective isometry between them that preserves the measures, see Definition 2.4 in [1]. On the other hand, the Gromov-weak topology is defined on weak equivalence classes of metric measure spaces: two spaces are weakly equivalent if their supports are strongly equivalent in the above sense, see Definition 2.1 in [19]. The Gromov–Hausdorff-weak topology is an adaptation of the GHP topology to weak equivalence classes; it is less standard, but more convenient to use when working with moments and Gromov-weak convergence. In order to obtain a GHP convergence, we will first prove the Gromov-weak convergence, then reinforce it to a Gromov–Hausdorff-weak one, and finally use the fact that the metric measure spaces considered here have full support to deduce the GHP convergence.
@@ 3.2.1 Ultrametric spaces and reduced processes
In this section we construct the reduced process associated with a general ultrametric space and prove that the function mapping an ultrametric space to its reduced process is continuous in the GHP topology.
@@
Suppose that $\big([U_{n},d_{n},\nu_{n}],\,n\geq 1\big)$ is a sequence of compact ultrametric measure spaces converging in the Gromov–Hausdorff–Prohorov topology to $U,d,\nu]$. Then
@@
This shows that there are at least $L_{t}(U)$ open balls of $(U_{n},d_{n})$ of radius $t-2\varepsilon$, that is,
@@
If the process $(L_{t}(U),t>0)$ only makes jumps of size $1$, the Skorohod convergence follows for instance from [28, Chapter VI, Theorem 2.15]. ∎
@@ 3.2.2 From Gromov-weak to GHP convergence
Let $\big([U_{n},d_{n},\nu_{n}],\,n\geq 1\big)$ be a sequence of random compact ultrametric measure spaces such that $\operatorname{supp}\nu_{n}=U_{n}$. Suppose that it converges to a limit $U,d,\nu]$ in distribution for the Gromov-weak topology. If, for all $t>0$, the collection of random variables
@@
is tight, then the sequence also converges in distribution for the Gromov–Hausdorff–Prohorov topology.
@@
In this result, $([U_{n},d_{n},\nu_{n}])_{n}$ should be interpreted as a random sequence of strong isometry classes such that the corresponding sequence of weak isometry classes converges in distribution in the Gromov-weak topology to $U,d,\nu]$. The limit of $([U_{n},d_{n},\nu_{n}])_{n}$ in the GHP topology is the unique strong isometry class with full support whose weak isometry class is $U,d,\nu]$.
@@
in the GHP topology. Since $U_{n},d_{n},\nu_{n}]$ has full support, $\iota([U_{n},d_{n},\nu_{n}])=[U_{n},d_{n},\nu_{n}]$ and the result is proved. ∎
@@ 4 The environmental coalescent point process
Fix some height $t>0$ and consider a Poisson point process $P$ on $0,\infty)\times(0,t]$ with intensity measure $ds\otimes\nu(dx)$, for some measure $\nu$ on $(0,t]$ such that
@@
see Figure 1 for an illustration. Consider a positive real $m$ and an independent exponentially distributed random variable $Z$ with mean $m$. The CPP at height $t$ corresponding to $\nu$ and $m$ is the random metric measure space $(0,Z),d_{\nu},\operatorname{Leb}]$. Its distribution will be denoted by $\operatorname{CPP}_{t}(\nu,m)$. Note that the Brownian CPP of Section 2.3 has distribution $\operatorname{CPP}_{1}(\tfrac{dx}{x^{2}},1)$.
@@
In the definition of a metric measure space of Section 3.1.1, we require the metric space to be complete and the measure to be defined on the corresponding Borel $\sigma$-field. Almost surely $(0,Z),d_{\nu}]$ is not complete, and we always work with its completion without further mention. The completion can be obtained explicitly by adding a countable number of points as described in [42]. Moreover, the Lebesgue measure $\operatorname{Leb}$ is formally defined on the usual $\sigma$-field on the interval $(0,Z)$, which might be different from the Borel $\sigma$-field of $(0,Z),d_{\nu}]$. The two $\sigma$-fields actually coincide. This issue has been addressed in details in [14]. For instance, [14, Lemma 1.4] shows that, if $\mathscr{I}$ is the usual Borel $\sigma$-field on $(0,Z)$ and $\mathscr{B}$ that induced by $d_{\nu}$, Point (iii) of [14, Definition 1.1] is fulfilled which entails that $\mathscr{I}\subseteq\mathscr{B}$ and
$\{y\in(0,Z):d_{\nu}(x,y)<t\}\in\mathscr{I},\quad t>0,\>x\in(0,Y).$
Since the topology induced by $d_{\nu}$ is separable, this shows that $\mathscr{I}=\mathscr{B}$. The Lebesgue measure is then implicitly extended to the completion of $(0,Z),d_{\nu}]$ by giving null mass to the countable number of points that were added.
The distribution of the rescaled genealogy of a nearly critical BPVE at time $tN$ will converge to $\operatorname{CPP}_{t}(\nu_{e},\bar{\rho}_{t})$, with $\bar{\rho}_{t}$ as in (16) and
@@
Let $(0,Z_{B}),d_{B},\operatorname{Leb}]$ be the Brownian CPP with distribution $\operatorname{CPP}_{1}\left(\tfrac{dx}{x^{2}},1\right)$. Recall the expression of $F$,
@@
and let $F^{-1}$ be its right-continuous inverse. Then $(0,Z_{B}),F^{-1}(d_{B}),\bar{\rho}_{t}\operatorname{Leb}]$ is distributed as $\operatorname{CPP}_{t}(\nu_{e},\bar{\rho}_{t})$, with $\nu_{e}$ defined in (30).
@@
Thus, $(0,Z_{B}),F^{-1}(d_{B}),\bar{\rho}_{t}\mbox{Leb}]$ is the CPP at height $t$ corresponding to the measure $\bar{\rho}_{t}\nu_{e}$ and to the random population size $Z_{B}$. By a simple scaling argument, it is not hard to see that the distribution of $(0,Z_{B}),F^{-1}(d_{B}),\bar{\rho}_{t}\operatorname{Leb}]$ is the same as that of $(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]$, the CPP constructed from $\nu_{e}$ and $Z_{e}$. (Note that the total mass needs to be scaled so that the isometry between $(0,Z_{B}),F^{-1}(d_{B})]$ and $(0,Z_{e}),d_{\nu_{e}}]$ is measure-preserving.) ∎
@@ 4.1 Moments of the CPP
For $k\geq 1$, let $(H_{1},\dots,H_{k-1})$ be independent and with cumulative distribution function $F$ defined in (16). As in (19), define an array
$\forall i\leq j,\quad H_{i,j}=H_{j,i}=\max\{H_{i},\dots,H_{j-1}\}.$ (31)
The moments of the environmental CPP can be computed from the law of this array as follows.
@@
Let $(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]$ be the environmental CPP at time $t$. Then, for any bounded function $\phi$ with corresponding polynomial $\Phi$,
@@
where $\left((H_{i,j})_{i,j}\right)$ is given by (31) and $\sigma$ is an independent and uniform permutation of $[k]$.
@@
$\forall s\leq t,\quad\mathbb{P}(H\leq s)=\frac{1/\bar{\rho}_{t}}{\nu_{e}((s,t])+1/\bar{\rho}_{t}}.$
This follows from the simple computation
$(\bar{\rho}_{t}\nu_{e}((s,t])+1)^{-1}=\frac{1}{2\bar{\rho}_{t}}\int_{[t-s,t]}e^{X_{t}-X_{u}}\sigma^{2}(du).\qed$
@@ 4.2 Sampling from the CPP
This problem would be much easier if the points were sampled according to an independent Poisson point process with intensity $\theta/\bar{\rho}_{t}dt$. The atoms of such a point process split $(0,Z_{e})$ into a geometric number of subintervals with success parameter $1/(1+\theta)$, and the lengths of these intervals are independent exponential random variables with mean $\bar{\rho}_{t}/(1+\theta)$. Thus, the distances between any two consecutive atoms are independent, and distributed as the maximum of $P$ over an interval with exponentially distributed length and mean $\bar{\rho}_{t}/(1+\theta)$. If $H^{\theta}$ denotes a random variable with the same distribution as this maximum, a direct computation shows that
@@
For $k\geq 1$, let $(U_{1},\dots,U_{k})$ be uniformly distributed on $(0,Z_{e})$. Then
$\mathbb{E}\big[\phi\big((d_{\nu_{e}}(U_{i},U_{j}))_{i,j}\big)\big]=k\int_{0}^{\infty}\frac{1}{(1+\theta)^{2}}\Big(\frac{\theta}{1+\theta}\Big)^{k-1}\mathbb{E}\big[\phi\big((H^{\theta}_{\sigma_{i},\sigma_{j}})_{i,j}\big)\big]d\theta,$ (34)
@@
$\forall x\geq 0,\quad e^{-x}=k\int_{0}^{\infty}\frac{1}{(1+\theta)^{2}}\Big(\frac{\theta}{1+\theta}\Big)^{k-1}f_{k+1,\theta+1}(x)d\theta,$
@@
$\int_{0}^{\infty}\frac{k}{(1+\theta)^{2}}\Big(\frac{\theta}{1+\theta}\Big)^{k-1}f_{k+1,\theta+1}(x)d\theta$ $=\int_{0}^{\infty}\frac{k}{(1+\theta)^{2}}\Big(\frac{\theta}{1+\theta}\Big)^{k-1}\frac{(\theta+1)^{k+1}x^{k}}{k!}e^{-x(\theta+1)}d\theta$ (35)
$=\int_{0}^{\infty}\frac{\theta^{k-1}x^{k}}{(k-1)!}e^{-x(\theta+1)}d\theta=e^{-x}.\qed$ (36)
@@
Let $(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]$ have distribution $\operatorname{CPP}_{t}(\nu_{e},\bar{\rho}_{t})$, let $(V_{1},\dots,V_{k})$ be i.i.d. uniform random variables on $(0,1)$, and define $U_{i}=V_{i}Z_{e}$, $i\leq k$. Since $Z_{e}$, $d_{\nu_{e}}$, and $(V_{i})_{i}$ are independent, using Lemma 4.5 we can write
$\mathbb{E}\big[\phi\big((d_{\nu_{e}}(U_{i},U_{j}))_{i,j}\big)\big]=k\int_{0}^{\infty}\frac{1}{(1+\theta)^{2}}\Big(\frac{\theta}{1+\theta}\Big)^{k-1}\mathbb{E}\big[\phi\big((d_{\nu_{e}}(V_{i}Z_{e},V_{j}Z_{e}))_{i,j}\big)\big]d\theta,$
@@
According to the previous discussion, the random variables $(H^{\theta}_{1},\dots,H^{\theta}_{k-1})$ are i.i.d. and distributed as the maximum of the point process $P$ over an interval with exponentially distributed length and mean $\bar{\rho}_{t}/(1+\theta)$. The distribution of this maximum is given by (33). By construction of the CPP distance,